OpenStudy (anonymous):

An investor invested a total of $800 in two mutual funds. One fund earned a 5% profit while the other earned a 3% profit. If the investor's total profit was $28, how much was invested in each mutual fund?

OpenStudy (anonymous):

So we have two variables \(a, b\) We know that \(a +b = 800\) and we know \(.05a + .03b =28 \) Two equations, two variables. Do you know how to solve a system of equations?

OpenStudy (anonymous):

is this the next step? .5(800-y)+.03y=28

OpenStudy (anonymous):

Yeah, seems you changed \(b\) into \(y\), but that's fine, lol. Keep going.

OpenStudy (anonymous):

450-.5y +.03y= 2800

OpenStudy (anonymous):

Wait, it should be \(.05\) not \(.5\)

OpenStudy (anonymous):

45-.05y+0.3=28.00

OpenStudy (anonymous):

450-0.5y+.03y=28.00

OpenStudy (anonymous):

450-.08y=$28.00

OpenStudy (anonymous):

\[.05(800-y) + .03y = 28 \Rightarrow .05(800) + .05y + .03y = 28\]

OpenStudy (anonymous):

so 40+.08y=28

OpenStudy (anonymous):

Yeah, so solve for \(y\), then plug that result into \(x+y=800\) to find \(x\). Then you will have the value for \(y\) and \(x\).

OpenStudy (anonymous):

Y=13 X=15 ?

OpenStudy (anonymous):

Oopos, I made a mistake. \[.05(800-y) + .03y = 28 \Rightarrow .05(800) - .05y + .03y = 28 \]

OpenStudy (anonymous):

So we have \[-.02y = 28-40 \Rightarrow y = 600 \]

OpenStudy (anonymous):

Sorry about that.

OpenStudy (anonymous):

x= 200

OpenStudy (anonymous):

Now \[x + 600 = 800\]

OpenStudy (anonymous):

Good

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Now what?

OpenStudy (anonymous):

5%= 600 3%=200

OpenStudy (anonymous):

cannot figure out what 5% and 3% would be

OpenStudy (anonymous):

cannot be true because if one is 600 and at 5% would be $30 and the other is $200 at 3% which would be $6.00 which equals $36. problem say $28 profit total

OpenStudy (anonymous):

No, you switched them around! It's 200 at 5% and 500 at 3%! Just check! It works.

hero (hero):

200 + 500 = 700, so that's not correct. Sorry buddy

OpenStudy (anonymous):

No, you switched them around! It's 200 at 5% and 600 at 3%! Just check! It works.

hero (hero):

If the user is confused about what the solution is, maybe you didn't explain it to him good enough.

OpenStudy (anonymous):

so 5% =$10 3% =$18 dont understand why at 3% would have more 5% how do you know which one is 600 and 200

OpenStudy (anonymous):

The way you know which belongs to which percent is based on the equation we had: \[.05x+.03y=28\] It is saying 5% belongs to x and 3% belongs to y.

OpenStudy (anonymous):

The reason why 3% profit had more total profit is because the initial investment (600) was larger than the other one (200).

hero (hero):

Also, an investor would probably invest more in a low risk account, whereas the 5% is likely to be more high risk investment, so you invest less.

OpenStudy (anonymous):

Okay thanks